# A Smooshed BMOCZ Zero Constellation for CFO Estimation Without Channel Coding

### v1

@@ I Introduction
Non-coherent communication schemes have emerged as a promising means to simplify receiver design and lower power consumption for ultra-massive connectivity in wireless networks [1]. Specifically, non-coherent schemes offer unique advantages in communication networks by eliminating the need for explicit CSI, which makes them appropriate for low data rate applications [2]. The lack of overhead in non-coherent schemes makes them especially beneficial in dynamic environments where CSI estimation may be impractical [3]. Since non-coherent schemes suffer from degraded performance relative to coherent ones, improving their reliability and spectral efficiency remains an active area of research.
@@
In [4], the authors propose a new non-coherent scheme called BMOCZ, where information bits are modulated onto the zeros of the baseband signal’s $z$-transform. The zeros are constrained to lie on one of two concentric circles in the complex plane, forming a zero constellation that the authors term Huffman BMOCZ. Prior studies have explored the optimal radius for BMOCZ and examined its integration into orthogonal frequency division multiplexing (OFDM) systems [5]. The primary advantage of BMOCZ is that the information zeros are preserved regardless of the channel realization, thereby making it non-coherent. This makes BMOCZ ideal for applications requiring ultra-reliable, low-latency communication, such as intermittent short-packet transmissions in Internet of Things (IoT) networks [6]. One notable use is over-the-air computation, a technique that struggles with channel-induced distortions and requires CSI estimation, which can be impractical in dynamic environments [7]. Additionally, the desirable auto-correlation properties of BMOCZ make it an ideal choice for integrated sensing and communication [8]. In general, BMOCZ offers specific advantages as a non-coherent scheme across a wide range of potential applications.
@@
When using BMOCZ for wireless communications, there are two notable impairments that can degrade performance: a time offset (TO) and a CFO [6]. A TO occurs whenever the start time of the transmitted signal misaligns with its actual arrival, which can lead to intersymbol interference (ISI). The authors in [6] exploit the auto-correlation properties of Huffman BMOCZ to estimate the TO and effective delay spread. The other impairment, a CFO, can significantly degrade performance. To address this problem in BMOCZ communication systems, the authors in [6] propose an affine cyclically permutable code (ACPC) combined with an oversampled direct zero-testing (DiZeT) decoder to estimate and correct the CFO. However, ACPC limits the code structure to CPCs. Therefore, to improve flexibility, alternative CFO correction methods for BMOCZ must be explored.
@@
In this study, we propose a new zero constellation called SBMOCZ, where the angular separation between adjacent zeros, excluding the first and last, is decreased. This creates a distinct gap in the zero constellation, which will rotate under a CFO. By identifying the gap location through the received polynomial, we can correct the CFO without channel coding. We note that our algorithm for estimating a CFO-induced rotation can be implemented via a single $N$-point discrete Fourier transform (DFT), making the implementation relatively efficient. Simulation results for BER and block error rate (BLER) in additive white Gaussian noise (AWGN) and fading channels show that SBMOCZ functions under a CFO without any coding, while having a modest performance loss compared to conventional BMOCZ in the absence of CFO. Furthermore, comparisons of coded SBMOCZ and BMOCZ using ACPC show that SBMOCZ achieves a $4$ dB BER gain in a fading channel while maintaining similar performance in other scenarios.
@@
Notation: The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$, respectively. The complex conjugate of $z=a+jb$ is expressed as $z^{\ast}=a-jb$. We denote the Euclidean norm of a vector $\boldsymbol{\rm{v}}\in\mathbb{C}^{N\times 1}$ as $||\boldsymbol{\rm{v}}||_{2}=\sqrt{\boldsymbol{\rm{v}}^{\text{H}}\boldsymbol{\rm{v}}}$. The circularly symmetric complex normal distribution with mean zero and variance $\sigma^{2}$ is expressed as $\mathcal{CN}(0,\sigma^{2})$. The uniform distribution on the interval $\left[a,b\right)$ is denoted by $\pazocal{U}_{[a,b)}$. We define $[N]=\{0,1,\dots,N-1\}$ to represent the set of the first $N$ non-negative integers. We represent the floor of a real number $a$ as $\lfloor a\rfloor$.
@@ II-A BMOCZ Fundamentals
For a BMOCZ constellation, the $k$th bit in a message $\boldsymbol{\rm{m}}=(m_{0},m_{1},\ldots,m_{K-1})\in\{0,1\}^{K}$ is mapped to the $k$th zero of a polynomial according to
$\alpha_{k}=\begin{cases}R_{k}\,e^{j\phi_{k}},&m_{k}=1\\ R_{k}^{-1}\,e^{j\phi_{k}},&m_{k}=0\end{cases}\>,$ (1)
for $k\in[K]$, $R_{k}>1$, and $\phi_{k}\in\left[0,2\pi\right)$. By the fundamental theorem of algebra, the zeros $\boldsymbol{\alpha}=(\alpha_{0},\alpha_{1},\dots,\alpha_{K-1})\in\mathbb{C}^{K}$ uniquely define the $K$th-degree polynomial
$X(z)=\sum_{k=0}^{K}x_{k}z^{k}=x_{K}\prod_{k=0}^{K-1}(z-\alpha_{k})\>,$ (2)
where $x_{K}\neq 0$. The polynomial coefficients, given by $\boldsymbol{\rm{x}}=(x_{0},x_{1},\dots,x_{K})\in\mathbb{C}^{K+1}$, are normalized such that their squared $L_{2}$-norm satisfies $||\boldsymbol{\rm{x}}||_{2}^{2}=K+1$. Let $W(z)=\sum_{n=0}^{N_{\rm z}-1}w_{n}z^{n}$ and $H(z)=\sum_{l=0}^{L-1}h_{l}z^{l}$ denote the $z$-domain representations of the noise sequence $\boldsymbol{\rm{w}}=(w_{0},w_{1},\dots,w_{N_{\rm z}-1})\in\mathbb{C}^{N_{\rm z}}$ and the $L$-tap channel impulse response $\boldsymbol{\rm{h}}=(h_{0},h_{1},\dots,h_{L-1})\in\mathbb{C}^{L}$, respectively. Assuming transmission through an LTI channel and applying the convolution theorem for $N_{\rm z}=K+L$, the received sequence can be expressed in the $z$-domain as
@@
where $\boldsymbol{\rm{y}}=(y_{0},y_{1},\dots,y_{N_{\rm z}-1})\in\mathbb{C}^{N}$ is the vector containing the coefficients of $Y(z)$. Although the presence of $H(z)$ introduces $L-1$ additional zeros to $Y(z)$, it does not alter the zeros corresponding to the message in $X(z)$. This property allows BMOCZ to operate without CSI, hence making it non-coherent. In this study, we assume a flat-fading channel where $L=1$, which means $N_{\rm z}=K+1$. This can achieved, for instance, through a time-frequency mapping of the BMOCZ polynomial coefficients in OFDM [5], which ensures $X(z)$ and $Y(z)$ have the same number of zeros.
The authors in [4] introduce Huffman BMOCZ, where all zeros are positioned on one of two concentric circles. This results in the zero mapping rule given in (1), with $R_{k}=r_{\text{hb}}$ and $\phi_{k}\triangleq 2\pi k/K$. In this scheme, $X(z)$ is called a Huffman polynomial, since the coefficients form a Huffman sequence for any combination of zeros [9, 10]. Huffman polynomials are well-conditioned, meaning that small variations in the coefficients lead to slight changes in the zeros, making them ideal for communication systems. The authors in [4] derive a simple decoding rule for BMOCZ called DiZeT, which evaluates $Y(z)$ at each conjugate-reciprocal zero pair $\mathscr{Z}_{k}=\{\alpha_{k},1/\alpha_{k}^{\ast}\}$. Using this method, the $k$th detected bit is given by
$\hat{m}_{k}=\begin{cases}1,&|Y(R_{k}\,e^{j\phi_{k}})|<R_{k}^{N_{\rm z}-1}|Y(R_{k}^{-1}\,e^{j\phi_{k}})|\\ 0,&\text{otherwise}\end{cases}\>.$ (4)
@@
According to [4], radial zero separation has a stronger influence on BER than angular separation. Therefore, the parameter $\lambda$ is introduced as a trade-off factor to balance radial versus angular zero separation.
@@ II-B CFO Impairment
A CFO occurs due to frequency mismatches between the transmitter and receiver oscillators. The presence of a CFO is a significant concern because it degrades the system performance. Let $\psi\in\left[0,2\pi\right)$ denote the amount of phase rotation caused by the CFO. [^1] Although this range of CFO-induced rotation is unrealistic in practice, and could disrupt the subcarrier orthogonality in OFDM, we consider it here for consistency with [6]. Due to this impairment, $Y(z)$ experiences a transformation in the $z$-domain
$\tilde{Y}(z)=\sum_{n=0}^{N_{\rm z}-1}y_{n}e^{j\psi n}z^{n}=Y(e^{j\psi}z)\>.$ (6)
Hence, all the zeros of $Y(z)$ rotate clockwise by the angle $\psi$. For a Huffman BMOCZ constellation with small $K$, this issue is less significant. However, for moderate to large values of $K$, even a small rotation causes decoding failure. Huffman BMOCZ, in the absence of channel coding, can only correct a fractional CFO due to the constant phase separation $\Delta{\phi}=2\pi/K$ between adjacent zeros and the uniform radii. Therefore, an angular rotation $\psi$ is only unique modulo $\Delta{\phi}$. Decomposing $\psi$ into a fractional component $\psi_{0}$ and an integer multiple of $\Delta{\phi}$, we obtain
$\psi=\psi_{0}+m\Delta{\phi}\;\;\text{where}\;\;\psi_{0}\in\left[0,\Delta{\phi}\right)\>,$ (7)
for any $m\in[K]$. Hence, only the fractional component $\psi_{0}$, relative to the angular separation $m\Delta{\phi}$, is detectable with uncoded Huffman BMOCZ. The authors in [6] propose using an ACPC combined with an oversampled DiZeT decoder to estimate and correct the total angular rotation. Within this scheme, the oversampled DiZeT decoder estimates the fractional component $\psi_{0}$, while the ACPC identifies the cyclic shift $m\Delta{\phi}$.
We highlight some limitations associated with ACPC. First, we note that the complexity of the code construction for ACPC increases when the code length is not a Mersenne prime, i.e., a prime number of the form $12^{K}-$ 1 for prime $K$. Furthermore, the approach requires cyclic codes, which confines the range of available coding structures. For instance, integrating BMOCZ with low-density parity check (LDPC) or polar codes is impractical within the ACPC framework. Finally, the approach introduces extra computational overhead in the decoding of the chosen cyclic code, which is required for CFO correction. For a detailed overview of ACPC, see the discussions in [6]. In this work, to retain flexibility, we propose a new BMOCZ constellation to address CFO-induced rotation without requiring any channel coding.
@@ III Methodology
The following subsections introduce a method to estimate and correct the CFO by using a smooshed BMOCZ zero constellation.
@@ III-A Smooshed Zero Constellation
We introduce a smooshed BMOCZ zero constellation where the phase difference between adjacent zeros, excluding the first and last, is given by $\Delta{\phi}=(2\pi-\zeta)/K$. The parameter $\zeta\in\left[0,2\pi\right)$, referred to as the smooshing factor, compresses the phase separation, which creates a larger gap between $\alpha_{0}$ and $\alpha_{K-1}$. The evaluation $|X(z)|$ along the unit circle is maximized closest to the center of the gap, since the density of zeros is lower in that region. As the constellation rotates, the gap shifts, which causes the maximum to move accordingly. Therefore, by evaluating the rotated polynomial $\tilde{Y}(z)$ at $z=e^{j\theta}$ for $\theta\in[0,2\pi)$, we can estimate $\psi$ from the $\theta$ that maximizes $|\tilde{Y}(e^{j\theta})|$. A detailed explanation for choice of the unit circle, along with a discussion of the algorithm for CFO estimation and correction, is presented in Section III-B.
@@
$\phi_{k}\triangleq\frac{(2\pi-\zeta)k}{K}+\frac{2\pi+\zeta\left(K-1\right)}{2K}\>.$ (8)
The choice of gap size affects the reliability of the CFO estimate and the displacement of zeros under noise. A small $\zeta$ results in poor CFO correction capabilities, while an excessively large $\zeta$ leads to significant zero perturbation under noise. Consequently, tuning the smooshing factor is crucial to obtain optimal performance. The choice of $\zeta$ also affects the minimum pairwise separation of the zeros, and hence the selection of radius. We modify (5) and obtain
$r_{\text{sb}}\triangleq\sqrt{1+2\lambda\sin{\left(\frac{2\pi-\zeta}{2K}\right)}}\>.$ (9)
@@
When $\zeta=0$, observe that (8) and (9) reduce to Huffman BMOCZ with a $\pi/K$ counterclockwise rotation of the zeros. In this study, for simplicity, we treat SBMOCZ with $\zeta=0$ and Huffman BMOCZ interchangeably. Fig. 1 illustrates the zero constellations for Huffman BMOCZ and SBMOCZ with $K=16$, where $\zeta=0.5$ for SBMOCZ. [^2] While $\zeta$ is typically much smaller in practice, we increase it here to emphasize the zero smooshing effect. The radii are computed by setting $\lambda=0.5$ and applying (5) and (9) to derive $R_{k}$ for Huffman BMOCZ and SBMOCZ, respectively.
@@
(b) SBMOCZ with $\zeta=0.5$.
@@
(a) Huffman BMOCZ
(b) SBMOCZ with $\zeta=3{\times}10^{-2}$
(c) SBMOCZ with $\zeta=5{\times}10^{-2}$

### v3

@@ I Introduction
Non-coherent communication has emerged as a promising means to simplify receiver design and lower power consumption for ultra-massive connectivity in wireless networks[1]. In particular, non-coherent communication offers unique advantages in wireless networks by eliminating the need for explicit CSI, which makes it suitable for low data rate applications[2]. The lack of overhead in non-coherent communication makes it especially beneficial in dynamic environments where CSI estimation may be impractical[3]. Since non-coherent communication suffers from degraded performance compared to coherent communication, improving its reliability and spectral efficiency remains an active area of research.
@@
In [4], the authors propose a novel non-coherent modulation scheme called BMOCZ, where the information bits are mapped to the zeros of the baseband signal’s $z$-transform. The information zeros are constrained to lie on one of two concentric circles in the complex plane, forming a zero constellation which the authors term Huffman BMOCZ. Previous studies have explored the optimal radius for BMOCZ, and examined its integration into orthogonal frequency division multiplexing (OFDM)[5]. The primary advantage of BMOCZ is that the information zeros are preserved at the receiver regardless of the channel realization, thereby making it non-coherent. This makes BMOCZ ideal for applications requiring ultra-reliable, low-latency communication, such as intermittent short-packet transmissions in Internet of Things (IoT) networks[6]. One notable use is over-the-air computation, a technique that struggles with channel-induced distortions and requires CSI estimation, which can be impractical in dynamic environments [7]. Moreover, the desirable auto-correlation properties of BMOCZ make it an ideal choice for integrated sensing and communication[8]. In general, BMOCZ offers specific advantages as a non-coherent modulation scheme across a wide range of potential applications.
@@
When utilizing BMOCZ for wireless communication, there are two notable impairments that can degrade the error rate performance: a time offset (TO) and CFO [6]. A TO occurs whenever the start time of the transmitted signal misaligns with its actual arrival, leading to intersymbol interference (ISI). In [6], the authors exploit the auto-correlation properties of Huffman BMOCZ to estimate and correct the TO. The CFO can also significantly degrade the error rate performance. To address the CFO in BMOCZ-based communication, the authors in [6] introduce an affine cyclically permutable code (ACPC) combined with an oversampled direct zero-testing (DiZeT) decoder to estimate and correct the CFO. However, the ACPC limits the code structure to CPCs. Therefore, to improve flexibility, alternative CFO correction methods for BMOCZ must be explored.
@@
In this study, we propose a new smooshed zero constellation called SBMOCZ, where the angular separation between adjacent zeros, excluding the first and last, is decreased. This creates a distinct gap in the zero constellation, which will rotate under a CFO. By identifying the gap location in the received polynomial, we can correct the CFO without channel coding. We highlight that our algorithm for estimating a CFO-induced rotation can be implemented via a single $N$-point discrete Fourier transform (DFT), making the implementation relatively efficient. The simulation results for BER and block error rate (BLER) in additive white Gaussian noise (AWGN) and fading channels demonstrate that SBMOCZ functions under a CFO without any channel coding, while incurring a modest performance loss compared to Huffman BMOCZ without a CFO. Furthermore, comparisons of coded SBMOCZ and Huffman BMOCZ with an ACPC show that SBMOCZ achieves a $4$ dB BER gain in the fading channel, while maintaining similar performance across other scenarios.
@@
Notation: The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$, respectively. We define $\mathbb{Z}^{N}_{2}$ as the set of all binary vectors of length $N$. The imaginary unit is given by ${\rm j}=\sqrt{-1}$, and Euler’s number is represented with ${\rm e}$. The complex conjugate of $z=a+{\rm j}b$ is expressed as $z^{\ast}=a-{\rm j}b$. We denote the Euclidean norm of a vector $\boldsymbol{\rm{v}}\in\mathbb{C}^{N\times 1}$ as $||\boldsymbol{\rm{v}}||_{2}=\sqrt{\boldsymbol{\rm{v}}^{\text{H}}\boldsymbol{\rm{v}}}$. The circularly symmetric complex normal distribution with mean zero and variance $\sigma^{2}$ is expressed as $\mathcal{CN}(0,\sigma^{2})$. The uniform distribution on the interval $\left[a,b\right)$ is denoted by $\pazocal{U}_{[a,b)}$. We define $[N]=\{0,1,\dots,N-1\}$ to represent the set of the first $N$ non-negative integers. Finally, we express the floor of a real number $a$ as $\lfloor a\rfloor$.
@@ II-A BMOCZ Fundamentals
[…]
$\alpha_{k}=\begin{cases}R_{k}\,{\rm e}^{{\rm j}\phi_{k}},&m_{k}=1\\ R_{k}^{-1}\,{\rm e}^{{\rm j}\phi_{k}},&m_{k}=0\end{cases}\>,\ \ k\in[K]\;,$ (1)
with $R_{k}>1$, and $\phi_{k}\in\left[0,2\pi\right)$. By the fundamental theorem of algebra, the zeros $\boldsymbol{\alpha}=[\alpha_{0},\alpha_{1},\dots,\alpha_{K-1}]\in\mathbb{C}^{K}$ uniquely define the $K$th degree polynomial
$X(z)=\sum_{k=0}^{K}x_{k}z^{k}=x_{K}\prod_{k=0}^{K-1}(z-\alpha_{k})\;,$ (2)
where $x_{K}\neq 0$. The polynomial coefficients, given by $\boldsymbol{\rm{x}}=[x_{0},x_{1},\dots,x_{K}]\in\mathbb{C}^{K+1}$, are normalized such that $||\boldsymbol{\rm{x}}||_{2}^{2}=K+1$. Let $W(z)=\sum_{n=0}^{N_{\rm z}-1}w_{n}z^{n}$ and $H(z)=\sum_{l=0}^{L-1}h_{l}z^{l}$ indicate the $z$-domain representations for the noise sequence $\boldsymbol{\rm{w}}=[w_{0},w_{1},\dots,w_{N_{\rm z}-1}]\in\mathbb{C}^{N_{\rm z}}$ and the $L$-tap channel impulse response $\boldsymbol{\rm{h}}=[h_{0},h_{1},\dots,h_{L-1}]\in\mathbb{C}^{L}$, respectively. Assuming transmission through an LTI channel and applying the convolution theorem for $N_{\rm z}=K+L$, the received sequence can be expressed in the $z$-domain as
@@
where $\boldsymbol{\rm{y}}=[y_{0},y_{1},\dots,y_{N_{\rm z}-1}]\in\mathbb{C}^{N_{\rm z}}$ is the vector containing the coefficients of $Y(z)$. Although the presence of $H(z)$ introduces $L-1$ additional zeros to $Y(z)$, it does not alter the zeros corresponding to the message in $X(z)$. This property allows BMOCZ to operate without CSI, thereby making it non-coherent. In this study, we assume a flat-fading channel where $L=1$, which means $N_{\rm z}=K+1$. This can be achieved, for example, through a time-frequency mapping of BMOCZ polynomial coefficients in OFDM, which ensures $X(z)$ and $Y(z)$ have the same number of zeros [5].
In [4], the authors introduce Huffman BMOCZ, where the information zeros are positioned on one of two concentric circles in the complex plane. This results in the zero mapping rule provided in (1), with $R_{k}=r_{\text{hb}}$ and $\phi_{k}\triangleq 2\pi k/K$. In this scheme, $X(z)$ is called a Huffman polynomial, since the coefficients form a Huffman sequence for any combination of information zeros [9, 10]. Huffman polynomials are well-conditioned, meaning that small variations in the coefficients lead to slight changes in the zeros, making them ideal for communication systems. In [4], the authors derive a simple decoding rule for BMOCZ called DiZeT, which evaluates $Y(z)$ at each conjugate-reciprocal zero pair $\mathscr{Z}_{k}\triangleq\{\alpha_{k},1/\alpha_{k}^{\ast}\}$. Using this method, the $k$th detected bit is given by
$\hat{m}_{k}=\begin{cases}1,&\left|Y(R_{k}\,{\rm e}^{{\rm j}\phi_{k}})\right|<R_{k}^{N_{\rm z}-1}\left|Y(R_{k}^{-1}\,{\rm e}^{{\rm j}\phi_{k}})\right|\\ 0,&\text{otherwise}\end{cases}\;.$ (4)
@@
According to [4], radial zero separation has a stronger influence on BER than angular separation. Therefore, the parameter $\lambda\in\left(0,1\right]$ is introduced as a trade-off factor to balance radial versus angular zero separation.
@@ II-B CFO Impairment
A CFO occurs due to frequency mismatches between the transmitter and receiver oscillators. The presence of a CFO in Huffman BMOCZ is a significant concern because it degrades the system performance. Let $\psi\in\left[0,2\pi\right)$ denote the amount of phase rotation caused by the CFO. [^1] Although this range of CFO-induced rotation is unrealistic in practice, and could disrupt the subcarrier orthogonality in OFDM, we consider it here for consistency with [6]. Due to this impairment, $Y(z)$ experiences a transformation in the $z$-domain
$\tilde{Y}(z)=\sum_{n=0}^{N_{\rm z}-1}y_{n}{\rm e}^{{\rm j}\psi n}z^{n}=Y(z{\rm e}^{{\rm j}\psi})\;.$ (6)
Hence, the zeros of $Y(z)$ rotate clockwise by the angle $\psi$. For a Huffman BMOCZ constellation with small $K$, this issue is less significant. However, for moderate or large values of $K$, even a small rotation causes decoding failure. Huffman BMOCZ, in the absence of channel coding, can only correct a fractional CFO due to the constant phase separation $\Delta{\phi}=2\pi/K$ between adjacent zeros and the uniform radii. Therefore, an angular rotation $\psi$ is only unique modulo $\Delta{\phi}$. Decomposing $\psi$ into a fractional component $\psi_{0}$ and an integer multiple of $\Delta{\phi}$, we obtain
$\psi=\psi_{0}+m\Delta{\phi}\;\;\text{where}\;\;\psi_{0}\in\left[0,\Delta{\phi}\right)\;,$ (7)
for any $m\in[K]$. Therefore, only the fractional component $\psi_{0}$, relative to the angular separation $m\Delta{\phi}$, is detectable with uncoded Huffman BMOCZ. To address this issue, the authors in [6] propose an ACPC combined with an oversampled DiZeT decoder to estimate and correct the total zero rotation. Within this framework, the oversampled DiZeT decoder estimates the fractional component $\psi_{0}$, while the ACPC identifies the cyclic shift $m\Delta{\phi}$.
We highlight several limitations associated with the ACPC. First, we note that the complexity of the code construction for an ACPC increases when the code length is not a Mersenne prime, i.e., a prime number of the form $12^{K}-$ 1 for prime $K$. Furthermore, the approach requires using cyclic codes, which confines the range of available coding structures. For instance, integrating BMOCZ with low-density parity check (LDPC) or polar codes is not practical within the ACPC framework. Finally, the approach introduces extra computational overhead in the decoding of the chosen cyclic code, which is required to correct the zero rotation. For a detailed overview of ACPC, we direct the reader to [6]. In this work, to retain flexibility, we propose a new BMOCZ zero constellation to address CFO-induced rotation without any channel coding.
@@ III Methodology
The following subsections introduce a framework to estimate and correct CFO-induced rotation by using a smooshed BMOCZ zero constellation.
@@ III-A Smooshed Zero Constellation
We introduce a smooshed BMOCZ zero constellation where the phase difference between adjacent zeros, excluding the first and last, is given by $\Delta{\phi}=(2\pi-\zeta)/K$. The parameter $\zeta\in\left[0,2\pi\right)$, referred to as the smooshing factor, compresses the phase separation, which creates a larger gap between $\alpha_{0}$ and $\alpha_{K-1}$. The evaluation of $|X(z)|$ along the unit circle is maximized closest to the center of the gap, since the density of zeros is lower in that region. As the constellation rotates, the gap shifts, which causes the maximum to move accordingly. Therefore, by evaluating the rotated polynomial $\tilde{Y}(z)$ at $z={\rm e}^{-{\rm j}\theta}$ for $\theta\in[0,2\pi)$, we can estimate $\psi$ from the $\theta$ that maximizes $|\tilde{Y}({\rm e}^{-{\rm j}\theta})|$. A detailed explanation for choosing the unit circle, along with a discussion of the algorithm for CFO estimation and correction, is provided in Section III-B.
@@
$\phi_{k}\triangleq\frac{(2\pi-\zeta)k}{K}+\frac{2\pi+\zeta\left(K-1\right)}{2K}\;.$ (8)
The selection of $\zeta$ affects the reliability of CFO correction and the displacement of zeros under noise. In particular, a small $\zeta$ results in poor CFO correction capabilities, while an excessively large $\zeta$ leads to significant zero perturbation under noise. Therefore, careful selection of the smooshing factor is required to obtain good error rate performance. The choice of $\zeta$ also affects the minimum pairwise separation of the zeros, and hence the selection of radius. We modify (5) and obtain
$r_{\text{sb}}\triangleq\sqrt{1+2\lambda\sin{\left(\frac{2\pi-\zeta}{2K}\right)}}\;.$ (9)
@@
When $\zeta=0$, observe that (8) and (9) reduce to Huffman BMOCZ with a $\pi/K$ counterclockwise rotation of the zeros. In this study, for simplicity, we treat SBMOCZ with $\zeta=0$ and Huffman BMOCZ interchangeably. Fig. 1 illustrates the zero constellations for Huffman BMOCZ and SBMOCZ with $K=16$, where $\zeta=1/2$ for SBMOCZ. [^2] While $\zeta$ is typically much smaller in practice, we increase it here to emphasize the zero smooshing effect. The radii are computed by setting $\lambda=1/2$ and applying (5) and (9) to derive $R_{k}$ for Huffman BMOCZ and SBMOCZ, respectively.
@@
(b) SBMOCZ with $\zeta=1/2$.
@@
(a) Huffman BMOCZ.
(b) SBMOCZ with $\zeta=3/100$.
(c) SBMOCZ with $\zeta=1/20$.